Can you solve this (c-4)^2 = 5
Answer by RAY100(1637)

take sqrt both sides
.
(c-4) = sqrt 5 = 2.236
.
c=4+2.236 = 6.236
.
.
check,,,, { (6.236) -4 }^2 = 5
.
2.236^2 =5,,,,,ok
.